OpenStudy (aryana_maria2323):

Please help! Will FAN AND MEDAL! Which of the following functions grows the fastest as x grows without bound? a. f(x) = x2 b. g(x) = x2 + 5x c. h(x) = the square root of the quantity x raised to the 4th power plus 2 times x d. They all grow at the same rate.

OpenStudy (aryana_maria2323):

Can you please help? @Directrix @mathmale

OpenStudy (aryana_maria2323):

@Agl202

OpenStudy (aryana_maria2323):

Any clue? @Agl202

OpenStudy (agl202):

@Directrix @dmezzullo Perhaps we could use some help here :)

OpenStudy (aryana_maria2323):

Can you please help? @mathmale

OpenStudy (mathmale):

First, Aryana, a bit of housekeeping: a. f(x) = x2 b. g(x) = x2 + 5x c. h(x) = the square root of the quantity x raised to the 4th power plus 2 times x It's important to express exponentiation with the ^ symbol: f(x)=x^2; g(x)=x^2+5x, \[h(x)=\sqrt{x^4+2x}\]

OpenStudy (aryana_maria2323):

a. x^2 b. x^2+5x c. yes your equation is correct

OpenStudy (mathmale):

One way of doing this would be to find the derivative of each function, as the derivative represents the instantaneous rate of change. Then you could rank order the derivatives from smallest to largest and thus answer the question at hand. a) the derivative with respect to x of x^2 is 2x; Find b) the deriv of x^2+5x and c) the deriv. of sqrt(x^4+2x). Compare them, assuming that x increases at the same rate for all of them.

OpenStudy (aryana_maria2323):

?? I am sorry but I do not understand?

OpenStudy (aryana_maria2323):

@mathmale

OpenStudy (aryana_maria2323):

@mathmale

OpenStudy (mathmale):

"Which of the following functions grows the fastest as x grows without bound?" All of the functions are increasing on their domains. The question here is: WHICH function increases fastest with x? The measure of rate of change is the DERIVATIVE. That's why I'm asking you to differentiate each of the 3 given functions.

OpenStudy (mathmale):

Derivative = instantaneous rate of change. Extremely important concept!

OpenStudy (mathmale):

rate of change = how fast something is increasing or decreasing as the independent variable increases. e. g., 50 mph: the distance traveled increases by 50 miles for each hour on the road.

OpenStudy (aryana_maria2323):

Okay so how do I find that out? Do I have to plug something into x?

OpenStudy (aryana_maria2323):

@mathmale

OpenStudy (aryana_maria2323):

Can you please help? @agent0smith

OpenStudy (agent0smith):

Do you know how to find derivatives...?

OpenStudy (aryana_maria2323):

I believe the answer is a?

OpenStudy (agent0smith):

Otherwise you can kinda compare them by looking at the highest power of x in each option, since the highest powers will dominate as x grows. Highest power of x in a is x^2, in b it's also x^2, and in c, it's effectively \[\large \sqrt {x^4}\]which can be simplified...

OpenStudy (aryana_maria2323):

to x^2 so the answer is actually D. Since they all have the same power right?

OpenStudy (agent0smith):

Yes

OpenStudy (aryana_maria2323):

Thank you. Can you help me with two more?

OpenStudy (aryana_maria2323):

?? @agent0smith

OpenStudy (agent0smith):

No, I still have to leave, I only came back for a while

OpenStudy (mathmale):

Please post each new problem separately. I'm curious: Why did you not answer agent0smith's question, "Do you know how to find derivatives...?" He was trying to determine the level of explanation that would be appropriate here.